IF explained variation is 80 and R^2 =0.8, then total variation is

A. 60

B. 80

C.100

D. 120
Answer #1

Answer

we can find the value of total variation by using the following formula

SST = SSR/R

where SST is total variation and SSR is explained variation

setting the value of SSR = 80 and R 0.80

we get

SST = (80/0.80)= 100

So, total variation is 100

option C is correct.